What volume (in L) of a 5.69 M Ca(OH)2 solution is required to titrate 576 mL of 2.32 M HNO3?

Give your answer to 3 sig figs.

Ca(OH)2(aq) + 2 HNO3(aq) → Ca(NO3)2(aq) + 2 H2O(l)
Answer #1

Balanced chemical equation is:
2 HNO3 + Ca(OH)2 ---> Ca(NO3)2 + 2 H2O


Here:
M(HNO3)=2.32 M
M(Ca(OH)2)=5.69 M
V(HNO3)=576.0 mL

According to balanced reaction:
1*number of mol of HNO3 =2*number of mol of Ca(OH)2
1*M(HNO3)*V(HNO3) =2*M(Ca(OH)2)*V(Ca(OH)2)
1*2.32 M *576.0 mL = 2*5.69M *V(Ca(OH)2)
V(Ca(OH)2) = 117 mL
V(Ca(OH)2) = 0.117 L

Answer: 0.117 L